OpenStudy (anonymous):

can you describe multiplying and dividing signed numbers in the real word? ( in your own words please ) :)

OpenStudy (anonymous):

@calculusxy

OpenStudy (calculusxy):

Is this saying like to explain the rules of multiplying and dividing with different signs?

OpenStudy (anonymous):

yes

OpenStudy (calculusxy):

Multiplying \[(+) \times (+) = (+)\] \[(-) \times (-) = (+)\] \[(+) \times (-) = (-)\] \[(-) \times (+) = (-)\] Dividing \[(+) \div (+) = (+)\] \[(-) \div (-) = (+)\] \[(+) \div (-) = (-)\] \[(-) \div (+) = (-)\]

OpenStudy (anonymous):

but can you explain it in words

OpenStudy (calculusxy):

I think that you can explain it from here. Can't you?

OpenStudy (anonymous):

no :(

OpenStudy (calculusxy):

just translate the symbols into words... it's pretty simple.

OpenStudy (anonymous):

but i dont know how to do that..can you do it please